# Flow Matching Meets PDEs: A Unified Framework for Physics-Constrained Generation

### v2

@@ Abstract
Generative machine learning methods, such as diffusion models and flow matching, have shown great potential in modeling complex system behaviors and building efficient surrogate models. However, these methods typically learn the underlying physics implicitly from data. We propose Physics-Based Flow Matching (PBFM), a novel generative framework that explicitly embeds physical constraints, both PDE residuals and algebraic relations, into the flow matching objective. We also introduce temporal unrolling at training time that improves the accuracy of the final, noise-free sample prediction. Our method jointly minimizes the flow matching loss and the physics-based residual loss without requiring hyperparameter tuning of their relative weights. Additionally, we analyze the role of the minimum noise level, $\sigma_{\min}$, in the context of physical constraints and evaluate a stochastic sampling strategy that helps to reduce physical residuals. Through extensive benchmarks on three representative PDE problems, we show that our approach yields up to an $8\times$ more accurate physical residuals compared to FM, while clearly outperforming existing algorithms in terms of distributional accuracy. PBFM thus provides a principled and efficient framework for surrogate modeling, uncertainty quantification, and accelerated simulation in physics and engineering applications. Source code and datasets are available at https://github.com/tum-pbs/PBFM.
@@ 1 Introduction
Partial differential equations (PDEs) constitute the foundational mathematical framework for describing the spatial and temporal evolution of physical systems (Evans, 2010). The numerical discretization of PDEs typically leads to high-dimensional systems, whose solution can become computationally prohibitive, particularly in the presence of nonlinearities or multiscale features (Haber et al., 2018; Valencia et al., 2025). To address these challenges, machine learning-based models have been proposed as an efficient alternative, capable of approximating the solution at a fraction of the computational cost (Chen et al., 2021; Fresca et al., 2021; Baldan et al., 2021; Brunton and Kutz, 2024). The framework of physics-informed neural networks (Raissi et al., 2019) rigorously embeds PDE constraints into the training process through automatic differentiation. However, this well-established approach is able to provide only a unique solution of the underlying PDE. Focusing only on deterministic systems is limiting if stochastic effects play a critical role, such as in the context of uncertainty quantification (Roy and Oberkampf, 2011; Abdar et al., 2021; Liu and Thuerey, 2024). Among generative modeling approaches, denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020; Nichol and Dhariwal, 2021), and implicit DDIM (Song et al., 2022), have been widely adopted across domains including image (Rombach et al., 2022), video (Ho et al., 2022), audio (Kong et al., 2021), and graph tasks (Chamberlain et al., 2021), demonstrating remarkable capabilities in reproducing complex data distributions. As an alternative to diffusion models, flow matching has emerged as a promising direction, providing a more direct and computationally efficient framework for generative modeling (Lipman et al., 2023). Furthermore, flow matching approach has shown the potential to achieve high-quality sample generation with a significantly reduced number of function evaluations (Esser et al., 2024).
The key challenge in integrating physical constraints into flow matching and diffusion-based generative models is resolving the fundamental conflict between two objectives: adhering to physical laws and accurately modeling the data distribution. The physical constraint term drives the model toward solutions that satisfy governing equations, typically favoring low-variance, deterministic outputs. In contrast, the generative objective promotes diversity and coverage of the data distribution, encouraging the model to increase variability. These goals are often in conflict: enforcing strict physical consistency can restrict the model’s capacity to represent the data distribution, while maximizing expressiveness may lead to physically invalid samples.
We propose Physics-Based Flow Matching (PBFM), a generative framework that addresses these shortcomings: our method retains the advantage of satisfying PDE constraints, while also enabling the reproduction of complex data distributions. It explicitly incorporates physical knowledge into the learning process by embedding it into the probability distributions. Our approach leverages temporal unrolling during training to refine the final state from which the physical residual is computed. This makes it possible to fully leverage the accurate gradient information from the embedded physical models. Importantly, the substantial gains in accuracy are achieved without negatively affecting inference performance. To the best of our knowledge, this represents the first integration of flow matching in a framework that embeds physical principles with high accuracy. Figure 1 provides a visual overview of the key training stages.
@@
Figure 1: The Physics-Based Flow Matching (PBFM) framework: During training, the sample $x_{t}$ at time $t$ is evolved to $t=1$ over $n$ time steps to compute the residual $\mathcal{R}(\widetilde{x}_{1})$. The flow matching loss $\mathcal{L}_{\text{FM}}$ and residual loss $\mathcal{L}_{\mathcal{R}}$ are combined in a conflict-free manner.
The key contributions of our work can be summarized as follows: (I) We introduce a novel framework that leverages flow matching to incorporate physical constraints, minimizing both PDE and algebraic residuals in a non-conflicting manner. Through numerical experiments across diverse test cases, we demonstrate that our approach outperforms state-of-the-art methods without the need to tune hyperparameters for balancing the loss terms. (II) We investigate the impact of temporal unrolling during training, showing that it enhances final predictions and reduces residual errors, without increasing inference time. (III) We highlight the effects of added Gaussian noise in flow matching when imposing a physical constraint. (IV) We compare the commonly used deterministic flow matching sampler with a stochastic alternative, revealing the latter’s superior performance in reproducing complex distributions.
@@ 2 Methodology
2.1 Physical residuals
A general time-dependent PDE in $n$-spatial dimensions can be expressed as $\mathbf{u}_{t}(\mathbf{x},t)=\mathcal{L}[\mathbf{u}(\mathbf{x},t)]+\mathbf{f}(\mathbf{x},t),\ \mathbf{x}\in\Omega\subseteq\mathbb{R}^{n}$, where $\mathbf{u}(\mathbf{x},t)$ denotes the unknown solution field, $\mathcal{L}$ is a spatial differential operator, and $\mathbf{f}(\mathbf{x},t)$ represents external forcing. To obtain numerical approximations, the spatial domain $\Omega$, along with its boundary $\partial\Omega$, is typically discretized, and continuous operators are replaced by their discrete counterparts (Karniadakis and Sherwin, 2005).
Any formulation that can be reduced to a residual for minimization can, in principle, be used to enforce the underlying physics. Residuals can be divided into three macro categories. In the case of steady-state PDEs, where the distribution of solution data typically stems from uncertainties in the underlying physics, the residuals can be directly computed from the governing equations as $\mathcal{R}(\mathcal{L}(\mathbf{u})+\mathbf{f})=0$. For time-dependent PDEs, where the learned solution evolves over time and consists of temporal snapshots, residuals are often formulated to enforce the conservation of physical quantities such as mass, momentum, or energy. Finally, algebraic constraints between different physical fields can also be incorporated as residuals, thereby further enforcing consistency with the underlying physics.
@@ 2.2 Physics-based flow matching
Integrating physical constraints into generative models often poses an inherently conflicting goal, leading to suboptimal distributional accuracy and trivially satisfied physical constraints. Existing, diffusion-based methods (Shu et al., 2023; Bastek et al., 2025) use the following optimization objective:
@@
where $\hat{r}$ are virtual observables (Rixner and Koutsourelakis, 2021) of the residual $\mathcal{R}(x_{1})$.
In the context of flow matching, the objective function to be minimized reduces to (Bastek et al., 2025):
@@
where $w_{\text{FM}}$ and $w_{\text{R}}$ are weights balancing the two losses. Specifically, increasing the weight of the physical residual term tends to degrade the quality of the learned distribution, while prioritizing the generative term undermines physical fidelity. Interestingly, multi-task learning has addressed the challenges of conflicting learning goals by formulating update directions based on gradients from the individual loss terms. We derive our method from the ConFIG method for computing conflict-free updates (Liu et al., 2025a), which in the context of physics-based flow matching yields:
@@
where $\mathbf{g}_{\text{FM}}$ and $\mathbf{g}_{\mathcal{R}}$ are the gradients for the flow matching loss $\mathcal{L}_{\text{FM}}$ and residual loss $\mathcal{L}_{\mathcal{R}}$, respectively. $\mathcal{O}(\mathbf{g}_{1},\mathbf{g}_{2})=\mathbf{g}_{2}-\frac{\mathbf{g}_{1}^{\top}\mathbf{g}_{2}}{|\mathbf{g}1|^{2}}\mathbf{g}_{1}$ is the orthogonality operator and $\mathcal{U}(\mathbf{g})=\mathbf{g}/|\mathbf{g}|$ is the unit vector operator. The obtained $\mathbf{g}_{\text{update}}$ satisfies $\mathbf{g}_{\text{update}}^{\top}\cdot\mathbf{g}_{\text{FM}}=\mathbf{g}_{\text{update}}^{\top}\cdot\mathbf{g}_{\mathcal{R}}>0$, which ensures that the gradient descent optimization along $\mathbf{g}_{\text{update}}$ always minimizes both losses simultaneously.
Crucially, this solves the gradient-level conflict between the objectives: rather than allowing one loss to dominate or oppose the other, this formulation ensures that both objectives contribute constructively to each optimization step adaptively. This prevents the optimizer from getting trapped in regions where one loss is minimized at the expense of the other, a common failure mode in standard weighted objectives. Consequently, our method enables the model to approximate a distribution close to pure flow matching while simultaneously reducing the enforced residual. Furthermore, we jointly optimize both terms by aligning their gradients into a unified update direction without relying on manual hyperparameter tuning. A comprehensive analysis is provided in Appendix E.
@@ 2.3 Improving physical and distributional accuracy
An important requirement for generative models in the context of physics applications is the necessity to produce high-accuracy samples that are on-par with traditional methods. We identify the reconstruction of the final, noise-free, sample during training as a crucial aspect for accurately evaluating the physical residual. The integration of the intermediate sample at time $t$ forward to the final time is performed with ODE integration (Lipman et al., 2024). Employing the predicted velocity field $u_{t}^{\theta}$ by integrating over a single time step of size 1t1-t. While trajectories are straight in theory, relying on just one integration step limits prediction accuracy, particularly when $t$ is close to zero. To improve the final prediction, we incorporate unrolling (Monga et al., 2021), by performing $n$ ODE integration steps of size (1t)/n(1-t)/n. The integration over multiple time steps has an averaging effect, which naturally promotes trajectories that are closer to the theoretically straight paths. Unrolling is applied via a curriculum, increasing the number of unrolling steps over the course of training. We additionally down-weigh less accurate predictions near $t=0$ by applying a scaling factor of $t^{p},\ p_{\text{opt}}=1$ in the residual loss computation. A detailed analysis of different power laws is provided in Appendix D. The more accurate prediction of $x_{1}$ substantially improves the evaluation of the physics residual in eq. 2, which in turn yields a better learning direction.
A final aspect of key importance in physics based flow matching is the value of $\sigma_{\text{min}}$, which denotes the amount of Gaussian noise added to the training data. In computer vision tasks, $\sigma_{\text{min}}$ is typically set to $10^{-3}$ (Lipman et al., 2023; Esser et al., 2024). However, the introduction of Gaussian noise perturbs the physical residuals, degrading the model’s performance. As a result, the amount of added noise establishes a lower bound on the residuals that the model can achieve, thereby directly influencing its ability to satisfy physical constraints. Thus, the $\sigma_{\text{min}}$ value has to be chosen carefully. We also found that $\sigma_{\text{min}}=0$ does not negatively affect the quality of the distribution and the underlying physics. Our explanation is that the numerical noise inherent to the physics residuals compensates for the lack of Gaussian mixing as required by flow matching theory.
@@
Inspired by natural image generation (Esser et al., 2024), we additionally explore sampling the time variable $t$ from a logit-normal distribution with zero mean and unit standard deviation during training, instead of the commonly used uniform distribution. This choice is motivated by the observation that flow matching tends to exhibit higher errors around $t=0.5$.
@@
Algorithm 1 details the resulting training procedure. The FM loss is not affected by the unrolling and is computed for accurately evaluating the physics at time $t$. The initial time is stored to weight the residual loss.
Algorithm 1 Training step for Physics Based Flow Matching
@@
At inference time, the final sample is typically obtained by evolving the initial noise, $\mathcal{N}(0,I)$ with ODE integration. This represents a deterministic sampler in which all the stochasticity is embedded in the initial noise sample (Gao et al., 2025). In contrast, the sampling process in DDPM is stochastic (Ho et al., 2020). Given that both diffusion models and flow matching can be seen generative modeling variants under arbitrary Markov processes (Holderrieth et al., 2025), we explore the use of a stochastic sampler within the physically based flow matching framework. The central idea is to evolve from time $t$ to $t=1$ and then return to $t+dt$ using a different noise sample. This step backwards in time with additional noise increases the stochasticity in the sampling process and improves distributional accuracy. The resulting procedure is outlined in Algorithm 2.
2.4 Ablation with a toy problem
To provide an intuition for the proposed improvements, we consider an ablation with a toy problem where the neural network outputs are the $x$ and $y$ coordinates of points on a circumference subject to the physical constraint of unit radius. Figure 2 shows the resulting point distribution along with the corresponding absolute physical error through mean estimation of the final sample. We compare with state-of-the-art models DM, as a DDPM representative (Song et al., 2022), and PIDM-ME, the best performing physics-informed algorithm (Bastek et al., 2025), and a PBFM-1 variant of our approach that omits unrolling. It is apparent that each extension yields a noticeable gain, and the final PBFM model exhibits a residual error that is 61.8 and 27.5 times lower than the DM and PIDM-ME baselines.

### v3

# Physics vs Distributions: Pareto Optimal Flow Matching with Physics Constraints

@@ Abstract
Physics-constrained generative modeling aims to produce high-dimensional samples that are both physically consistent and distributionally accurate, a task that remains challenging due to often conflicting optimization objectives. Recent advances in flow matching and diffusion models have enabled efficient generative modeling, but integrating physical constraints often degrades generative fidelity or requires costly inference-time corrections. Our work is the first to recognize the trade-off between distributional and physical accuracy. Based on the insight of inherently conflicting objectives, we introduce Physics-Based Flow Matching (PBFM) a method that enforces physical constraints at training time using conflict-free gradient updates and unrolling to mitigate Jensen’s gap. Our approach avoids manual loss balancing and enables simultaneous optimization of generative and physical objectives. As a consequence, physics constraints do not impede inference performance. We benchmark our method across three representative PDE benchmarks. PBFM achieves a Pareto-optimal trade-off, competitive inference speed, and generalizes to a wide range of physics-constrained generative tasks, providing a practical tool for scientific machine learning.
Code and datasets available at https://github.com/tum-pbs/PBFM.
@@ 1 Introduction
Partial differential equations (PDEs) provide the core mathematical framework for modeling the evolution of physical systems across space and time (Evans, 2010). However, discretizing PDEs often results in high-dimensional problems that are computationally prohibitive, especially for nonlinear or multiscale phenomena (Haber et al., 2018; Valencia et al., 2025). Recently, machine learning-based methods have emerged as efficient alternatives, enabling the approximation of PDE solutions with significantly reduced computational cost (Chen et al., 2021; Fresca et al., 2021; Baldan et al., 2021; Brunton & Kutz, 2024). Physics-informed neural networks (PINNs) (Raissi et al., 2019) embed PDE constraints directly into the training objective via automatic differentiation. However, PINNs yield a single deterministic solution, limiting their use in scenarios where stochasticity is crucial, such as uncertainty quantification (Roy & Oberkampf, 2011; Abdar et al., 2021; Liu & Thuerey, 2024). In contrast, generative models like denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020; Nichol & Dhariwal, 2021), their implicit variants (DDIMs) (Song et al., 2022), and flow matching (Lipman et al., 2023) have shown strong performance in capturing complex data distributions across domains, including images, videos, audio, and graphs (Rombach et al., 2022; Ho et al., 2022; Kong et al., 2021; Chamberlain et al., 2021). Flow matching, in particular, offers a conceptually simple and computationally efficient framework, achieving high-fidelity sample generation with fewer function evaluations (Esser et al., 2024).
Despite recent progress, integrating physical constraints into generative models remains challenging. Enforcing such constraints during training requires the model to jointly optimize for generative fidelity and physical consistency, often resulting in conflicting gradients that can compromise either distributional accuracy or physical validity (Krishnapriyan et al., 2021). Achieving a balanced trade-off between these objectives is difficult, as improvements in one can degrade the other. In this work, we revisit physics-constrained generative modeling from a Pareto perspective, explicitly targeting solutions that balance physical correctness and generative performance; we systematically analyze the trade-offs involved.
Additionally, the so-called Jensen’s gap exists: a fundamental discrepancy that arises when physical constraints, such as PDE residuals, are imposed on the posterior mean $\mathbb{E}[x_{1}|x_{t}]$ at intermediate noise levels, rather than directly on the final clean sample $x_{1}$ (Zhang & Zou, 2025). Alternatively, enforcing constraints only at inference time typically requires iterative sampling procedures to achieve physical consistency, which can be computationally expensive and undermine the efficiency of generative models. Moreover, enforcing nonlinear hard constraints necessitates differentiating through the physical residual, a process that is computationally expensive for high-dimensional and complex systems (Cheng et al., 2025).
We improve constrained flow matching with a series of established techniques such as unrolling, residual-based losses, stochastic sampling, and conflict-free gradient updates (Liu et al., 2025a), adapting them to address the unique challenges of physics-constrained generative modeling. Through the lens of the inherent physics-vs-distribution trade-off, we show that these modifications yield very substantial improvements over previous work, and enable training networks that yield distributional accuracy and satisfy physics constraints at the same time. Fig. 1 illustrates the overall training procedure and highlights the key components of PBFM method.
@@
Figure 1: Overview of the proposed Physics-Based Flow Matching (PBFM) approach. During training, the sample $x_{t}$ at time $t$ is evolved to $t=1$ over $n$ time steps to compute the residual $\mathcal{R}(\widetilde{x}_{1})$. The flow matching loss $\mathcal{L}_{\text{FM}}$ and residual loss $\mathcal{L}_{\mathcal{R}}$ are combined in a conflict-free manner.
Our main contributions are as follows. (I) We introduce a method that integrates physical constraints into flow matching, enabling the minimization of both PDE and algebraic residuals without manual balancing. (II) We show that unrolling during training effectively mitigates Jensen’s gap, leading to lower residual errors and improved final predictions without increasing inference cost. (III) We analyze the role of Gaussian noise in flow matching under physical constraints, and (IV) we conduct a comprehensive comparison between deterministic and stochastic flow matching samplers, demonstrating advantages of the latter. A central advantage of the proposed approach is that it is very easy to implement in existing flow matching pipelines, and, as we will demonstrate below, consistently achieves substantial improvements in terms of distributional and physical accuracy.
2 Related work
Numerous deep learning methods have been developed to address complex problems in physics and engineering (Morton et al., 2018; Wang et al., 2020; Sanchez-Gonzalez et al., 2020; Thuerey et al., 2020). More recently, there has been growing interest in bringing the foundation model paradigm to scientific machine learning. Efforts such as PDEformer (Ye et al., 2024), Poseidon (Herde et al., 2024), Aurora (Bodnar et al., 2024), and Unisolver (Zhou et al., 2025a) aim to build models with broad generalization capabilities across diverse physical systems. In parallel, specialized transformer architectures tailored for PDEs have also been proposed, including OFormer (Li et al., 2023), Transolver (Wu et al., 2024), Fengbo (Pepe et al., 2025), and PDE-Transformer (Holzschuh et al., 2025).
Focusing on diffusion models in the physics and engineering fields, representative applications include the generation and design of new molecules and drugs (Guo et al., 2024; Schneuing et al., 2024; Bose et al., 2024; Guastoni & Vinuesa, 2025), the simulation of particle trajectories in collider experiments (Mikuni et al., 2023), solving inverse PDE problems (Holzschuh et al., 2023), and the modeling of particle motion in turbulent flows (Li et al., 2024). Alongside these efforts, several works have sought to embed physical knowledge or constraints directly into the diffusion process to enhance model performance. For instance, Huang et al. (2024) presented an approach to solve PDEs from partial observations by filling in missing information using generative priors. Zhou et al. (2025b) proposed an approach that embeds priors into the diffusion process to satisfy energy and momentum conservation laws and PDE constraints. Rixner & Koutsourelakis (2021) formulated a probabilistic generative model enforcing physical constraints through virtual observables.
More aligned with our work, Shu et al. (2023) proposed an algorithm that conditions the diffusion process on the residual gradient both at training and inference time, with a focus on super-resolution and reconstruction tasks based on random measures of turbulent flows. Despite being one of the first works in this direction, they do not directly enforce the residual. Similarly, Jacobsen et al. (2025) introduced CoCoGen, a method that employs the governing PDEs during inference, while leaving the training procedure unchanged. The approach improves the physical residual but slows down inference, which is crucial in real applications. Other approaches that enforce physical constraints at inference time include FFM (Kerrigan et al., 2023), DiffusionPDE (Huang et al., 2024), D-Flow (Ben-Hamu et al., 2024), ECI (Cheng et al., 2025), and PCFM (Utkarsh et al., 2025). These methods require iterative sampling procedures to achieve physical consistency, which can be computationally expensive and can take more than $10\times$ longer than the standard diffusion process (Utkarsh et al., 2025). The last two methods are designed to strictly satisfy the constraints, but the ECI method is limited to simple non-overlapping constraints, while PCFM is applicable to arbitrary constraints but requires differentiating through the residual, which is computationally expensive. In contrast, Bastek et al. (2025) proposed physics-informed diffusion models (PIDM), which incorporate an additional loss term during training to minimize physical residuals. Zhang & Zou (2025) introduced physics-informed distillation of diffusion models (PIDDM), to potentially solve the Jensen’s gap issue by distillation that is fine-tuned to minimize the physical residual. However, this approach requires training two separate models and, like Bastek et al. (2025) does not resolve the inherent conflict between the generative and physical objectives. Other recent methods that adopt fine-tuning to reduce the physics constraints are PIRF (Yuan et al., 2025) and PCFT (Tauberschmidt et al., 2025). We provide a comparison of our method with existing physics-constrained generative models in Table 1.
Table 1: Comparison between our method and other physics-constrained generative models.
Method Physics at Hard Gradient Free Complex Balanced
Training Constraints Inference Constraints Hyperparams
FFM [34]
ProbConserv [24]
CoCoGen [32]
DiffusionPDE [31]
PIDM [9]
D-Flow [10]
ECI [16]
PCFM [64]
PIDDM [71]
PBFM (ours)
@@ 3 Methodology
3.1 Preliminaries
A general time-dependent PDE in $n$ spatial dimensions can be written as $\mathbf{u}_{t}(\mathbf{x},t)=\mathcal{L}[\mathbf{u}(\mathbf{x},t)]+\mathbf{f}(\mathbf{x},t),\quad\mathbf{x}\in\Omega\subseteq\mathbb{R}^{n},$ where $\mathbf{u}(\mathbf{x},t)$ is the solution field, $\mathcal{L}$ is a spatial differential operator, and $\mathbf{f}(\mathbf{x},t)$ denotes external forcing. For numerical approximation, the domain $\Omega$ and its boundary $\partial\Omega$ are discretized, and continuous operators are replaced by discrete analogues (Karniadakis & Sherwin, 2005).
Physics constraints can be enforced by minimizing a residual function, which is constructed drawing from a wide range of formulations. In generative modeling, the stochasticity of residuals can be categorized into three main types: (i) For steady-state PDEs, the solution distribution arises from uncertainties in the physical parameters, and the residual is computed directly from the governing equations, e.g., $\mathcal{R}(\mathcal{L}(\mathbf{u})+\mathbf{f})=0$. (ii) For time-dependent PDEs, the solution evolves over time and residuals are formulated to enforce conservation laws (such as mass, momentum, or energy) across temporal snapshots. (iii) Algebraic constraints can be used to impose additional consistency between physical fields, further enforcing the underlying physics of the system.
@@ 3.2 Physics-based flow matching
Integrating physical constraints into generative models introduces an inherent conflict: optimizing for physical fidelity often degrades distributional accuracy, and vice versa. Existing diffusion-based approaches (Shu et al., 2023; Bastek et al., 2025) typically employ a weighted objective:
@@
where $\hat{r}$ are virtual observables (Rixner & Koutsourelakis, 2021) of the residual $\mathcal{R}(x_{1})$. For flow matching, this reduces to (Bastek et al., 2025):
@@
where $w_{\text{FM}}$ and $w_{\mathcal{R}}$ are adjusted manually to address the potential conflicts between generative and physical losses. However, tuning these weights is challenging: increasing the residual term often harms generative quality, while prioritizing the generative term undermines physical consistency.
To address this, we leverage multi-task optimization techniques that resolve conflicting gradient directions. Specifically, we adopt the ConFIG method (Liu et al., 2025a) to compute conflict-free updates for physics-based flow matching:
@@
Figure 2: PBFM method leverages ConFIG to combine the unit vectors of the orthogonal components and scales the result using the projection length of the FM and the physical residual $\mathcal{R}$ gradients.
where $\mathbf{g}_{\text{FM}}$ and $\mathbf{g}_{\mathcal{R}}$ are the gradients of $\mathcal{L}_{\text{FM}}$ and $\mathcal{L}_{\mathcal{R}}$, respectively; $\mathcal{O}(\mathbf{g}_{1},\mathbf{g}_{2})=\mathbf{g}_{2}-\frac{\mathbf{g}_{1}^{\top}\mathbf{g}_{2}}{|\mathbf{g}_{1}|^{2}}\mathbf{g}_{1}$ is the orthogonality operator, and $\mathcal{U}(\mathbf{g})=\mathbf{g}/|\mathbf{g}|$ normalizes to unit length. The resulting update direction $\mathbf{g}_{\text{update}}$ guarantees simultaneous descent on both objectives: $\mathbf{g}_{\text{update}}^{\top}\mathbf{g}_{\text{FM}}>0$ and $\mathbf{g}_{\text{update}}^{\top}\mathbf{g}_{\mathcal{R}}>0$. Figure 2 represents the gradient composition.
This approach adaptively aligns the gradients, eliminating the need for manual loss weighting and preventing one objective from dominating optimization. As a result, the resulting updates yield high distributional accuracy while still enforcing physical constraints, consistently outperforming fixed-weighted objectives across a wide range of weights (see Appendix F for an analysis).
@@ 3.3 Improving physical and distributional accuracy
A key requirement for generative models in scientific applications is the ability to produce samples with high physical accuracy, close to traditional numerical methods. In physics-constrained generative modeling, a central challenge is the so-called Jensen’s gap which appears whenever a nonlinear map $f$ is applied to a random variable $Z$: in general $\mathbb{E}[f(Z)]\neq f(\mathbb{E}[Z])$. When physical constraints are imposed on the posterior mean $\mathbb{E}[x_{1}|x_{t}]$ at intermediate noise levels, rather than directly on the final clean sample $x_{1}$, a discrepancy arises that can degrade physical fidelity (Zhang & Zou, 2025). To mitigate this, we identify the reconstruction of the final, noise-free sample during training as crucial for accurately evaluating the physical residual. Specifically, we employ unrolling, i.e., integrating the intermediate sample at time $t$ forward to the final time $t=1$ using multiple ODE steps. This process mitigates Jensen’s gap by ensuring that the residual is evaluated on a more accurate prediction of $x_{1}$, rather than a single-step approximation. By unrolling over $n$ steps of size (1t)/n(1-t)/n, the integration better approximates the true trajectory. Unrolling is applied via a curriculum, gradually increasing the number of steps during training. Additionally, we down-weight less accurate predictions near $t=0$ by applying a scaling factor $t^{p}$ (with $p_{\text{opt}}=1$) to the residual loss, further improving the learning signal. Linear scaling is also principled, aligning with the linear noise schedule in flow matching and ensuring consistency between residual weighting and FM dynamics. An investigation of different power laws is provided in Appendix E. The improved prediction of $x_{1}$ substantially enhances the evaluation of the physics residual in Eq. 2, yielding better optimization directions and directly mitigating Jensen’s gap. This leads to lower residual errors and improved final predictions, without increasing inference cost. Nevertheless, unrolling increases memory consumption during training, as intermediate states must be stored for backpropagation. Table 6 details the additional training time and memory requirements for different numbers of unrolling steps.
Another important aspect in physics-based flow matching is the choice of $\sigma_{\text{min}}$, the amount of Gaussian noise added to training data adhering to the mixture of Gaussians theory. While computer vision tasks typically use $\sigma_{\text{min}}=10^{-3}$ (Lipman et al., 2023; Esser et al., 2024), excessive noise perturbs physical residuals and degrades performance. The value of $\sigma_{\text{min}}$ thus influences the minimum achievable residual error. We provide an analysis of different noise levels in Table 5. A practical guideline is that adding Gaussian noise of scale $\sigma_{\text{min}}$ induces a residual MSE $\approx\sigma_{\text{min}}^{2}$ in a perfect reconstruction setting, reducing to $\sigma_{\min}\lesssim\mathcal{R}_{\min}$.
@@
Inspired by natural image generation (Esser et al., 2024), we also sample the time variable $t$ from a logit-normal distribution (zero mean, unit standard deviation) during training, instead of a uniform distribution. This specifically targets regions where flow matching exhibits higher errors, typically around $t=0.5$.
@@
Algorithm 1 details the resulting training procedure, where the FM loss is computed at time $t$ and unrolling is used for accurate residual evaluation. The initial time is stored to weight the residual loss.
Algorithm 1 Training procedure for PBFM
@@
At inference time, the final sample is typically obtained by evolving the initial noise, $\mathcal{N}(0,I)$ with ODE integration. This represents a deterministic sampler in which all the stochasticity is embedded in the initial noise sample (Gao et al., 2025). In contrast, the sampling process in DDPM is stochastic (Ho et al., 2020). Given that both diffusion models and flow matching can be seen generative modeling variants under arbitrary Markov processes (Holderrieth et al., 2025), we explore the use of a stochastic sampler, similarly to what has been applied in ECI method (Cheng et al., 2025), within the physically based flow matching framework. The central idea is to evolve from time $t$ to $t=1$ and then return to $t+dt$ using a different noise sample. This step backwards in time with newly generated noise increases the stochasticity in the sampling process and improves distributional accuracy. The resulting procedure is outlined in Algorithm 2.